OpenStudy (nottim):

What is the limiting reagent and % yield of NH3 in N2+3H2--->2NH3

OpenStudy (nottim):

Given 14.02g of N2 and 3.15 g H2 and 14.5g NH3

Parth (parthkohli):

That's hot. Write down the moles and stuff.

OpenStudy (astrophysics):

n=m/M

Parth (parthkohli):

I suppose you're using exact molar masses. Ew. I've only memorised them to the nearest integer.

Parth (parthkohli):

Alright, so any ideas? Tim?

OpenStudy (nottim):

I'm sorry, I was helping someone else with this question, but she has NO idea.

OpenStudy (nottim):

but i got 0.9995 mol N2, 3.125 mol H2, 0.85138 mol

Parth (parthkohli):

Cool. What's the limiting reagent then?

OpenStudy (nottim):

Nitrogen. But my friend will see this answer.

Parth (parthkohli):

Cool, so proportionally, 3.125 mol. H2 will react with 3.125/3 mol. N2 = 1.04 mol. N2, but we don't have that much of nitrogen so yeah, that's the limiting reagent.

OpenStudy (nottim):

% yield, thats just multiplying h2 by 3, and nh3 by 2 right

Parth (parthkohli):

So like we should have gotten 0.95 * 2 = 1.90 moles of NH3, but we didn't even get that much

OpenStudy (nottim):

i multiply by them by their subscripts and mole count?